OpenStudy (anonymous):

(3,-4) (2,1) a. -5 b. -3/5 c. -1

OpenStudy (anonymous):

@satellite73 @mathstudent55 @Jasmineflvs @phi

OpenStudy (anonymous):

another "rise over run" right?

OpenStudy (anonymous):

\[\frac{-4-1}{3-2}\]

OpenStudy (anonymous):

@satellite73 yea :/

OpenStudy (anonymous):

but really you should do this in your head

OpenStudy (anonymous):

@satellite73 when I do that I divide what i got on the top to the bottom and thats the answer?

OpenStudy (anonymous):

from 2 to 3 is right two from 1 to -4 is down five right one, down five, slope is \(-5\)

OpenStudy (anonymous):

yeah,that is what you have to do find the top and bottom and then divide

OpenStudy (anonymous):

\[\frac{-4-1}{3-2}=\frac{-5}{1}=-5\]

OpenStudy (anonymous):

@satellite73 oh that is pretty easy! thank you

OpenStudy (anonymous):

yeah it is easy enough yw